Use the complement rule to compute probabilities

Question

A spinner contains the numbers 1 through 75 . What is the probability that the spinner will land on a number that is not between 16 and 25 , inclusive?
Answer #1

A spinner contains numbers 1 through 75

P( spinner will land on a number that is not between 16 and 25)

= 1 - P( spinner will land on a number that is between 16 and 25)

= 1 - 10/75

= 65/75

= 0.8667

P( the spinner will land on a number that is not between 16 and 25) = 0.8667